how would you graph and solve 3y>4x?
Answer by philline_palana(20)

3y>4x


3y=4x
if x = -1
3y=4(-1)
3y/3=-4/3
y=-4/3
Point 1: (-1, -4/3)

if x = -2
3y=4(-2)
3y/3=-8/3
y=-8/3
Point 2: (-2, -8/3)

if x = -3
3y=4(-3)
3y/3=-12/3
y=-4
Point 3: (-3, -4)

if x = 1
3y=4(1)
3y/3=4/3
y=4/3
Point 4: (1, 4/3)

if x = 2
3y=4(2)
3y/3=8/3
y=8/3
Point 5: (2, 8/3)

if x = 3
3y=4(3)
3y/3=12/3
y=4
Point 5: (3, 4)

connect the points.
Shade the upper part of the line because y>x